OpenStudy (anonymous):

What is the additive inverse of 4x? A. 4 – x B. –4x C. 1/4x D. x – 4

OpenStudy (anonymous):

4x + something = 0 Solve for something and you have your answer.

Directrix (directrix):

Hint: 4 + ? = 0

OpenStudy (anonymous):

i don't get it..

OpenStudy (anonymous):

Does 4x + (4 -x) = 0? Does 4x + (-4x) = 0? Does 4x + (1/4x) = 0? Does 4x + (x - 4) = 0? Only one of them is true.

Directrix (directrix):

Read the attached file and see if that helps. I believe it will.

OpenStudy (anonymous):

so is it 4-x

Directrix (directrix):

I need to know this answer first: 4 + ? = 0 Please post what ? =

OpenStudy (anonymous):

4+ -4

Directrix (directrix):

4+ -4 = 0. Yes. So -4 and +4 are additive inverses. Now for the big one. ============ 4x + (-4x) = what? Post what the value of what is.

Directrix (directrix):

Are 4x and -4x additive inverses? Do they add to 0? Yes or no.

Directrix (directrix):

@brittbrat213

OpenStudy (anonymous):

yes i think

Directrix (directrix):

Yes, they do. So, Does 4x + (-4x) = 0? The answer to the question is yes. (the correct option of those you posted)

OpenStudy (anonymous):

then the answer is b

Directrix (directrix):

Correct